OpenStudy (mary.rojas):

The system of equations represented by x+2y=2 and -4x+3y=25 intersects at (a,b). What does -2(a+b)=? a)2 b)4 c)12 d)24 E)30

OpenStudy (anonymous):

x+2y=2 -4x+3y=25 solve for y and x by elimination or substitution or graphing. then. you plug into that formula -2(x+y)

OpenStudy (mary.rojas):

whats substitution again... :)

OpenStudy (harsimran_hs4):

substitution is a method in which you find a variable in terms of other (from one equation) and then substitute it in another equation say in this ques x = 2-2y (from 1st eq) so put this value of x in second and solve for x then take this value of x and put it in any of the eq to get y

OpenStudy (mary.rojas):

i get y=3 and x=-4 so the final answer i get is 2.... so is the answer "A" which is 2 ?

OpenStudy (mary.rojas):

pleaseeeeeeee someone is the answer "A" which is 2

HanAkoSolo (jamierox4ev3r):

-2(a+b) = -2(3+(-4))= -2(3-4) = -6+8=2 2 is your answer , so that is correct. However, do A and B necessarily signify X and Y?

OpenStudy (mary.rojas):

okay you lost me at your question... "However, do A and B necessarily signify X and Y?"

OpenStudy (mary.rojas):

p.s- thank you for your reply !